• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 04:50:15 on February10,2009

    Entry number 260188

    keyword=Compton photon rate

    Looking at the photon rate we send to MCC, I am starting to think that the signal which clears the scaler might not be good and is clearing the scaler not for all helicity state. The rate stays at 0 if a run is started with detector off, gets crazy when detector is turned on, and stays non zero after turning the detector off. Playing with the gain, threshold and attenuation did not really help. I turned the beam diagnostic PMT and it seems the plots in spydiag also have access to another scaler for the central crystal, so rates seems to be of the order 500 Hz for 4uA. Looking with the oscilloscope I found also a rate of roughly 700 Hz.
    So I think I conclude that the beam is well tuned and the electronics has a problem and will have to be checked at next access. I pulled out the spydiag screen in font room so they can keep an eye on the rate. I will stay in the compton chicane if the FADC crew wants to take data on the GSO, if rates gets crazy I asked to be paged and will decide if it is safer to go straight but beam seems reasonably stable so I guess I will sleep well.


    FIGURE 1

    FIGURE 2